# Performance Estimation in Binary Classification Using Calibrated Confidence

### v1

@@ 3.2.3 Recall
The joint distribution of $X_{TP}$ and $X_{FN}$ is again just the product of the marginal distributions since $X_{TP}\perp\!\!\!\!\perp X_{FN}$. Thus, we can form the distribution of $X_{recall}$ by iterating over the joint and collecting all encountered values of the expression on the right side of the Equation (7) with their respective probabilities. A small shortcut is to realize that $X_{recall}=0$ whenever $X_{TP}=0$, and $X_{recall}=1$ whenever $X_{FN}=0$, with the exception of both $X_{TP}$ and $X_{FN}$ being zero, which means that recall is undefined. In most practical situations, the probability of this event is negligible, but a user should be prepared to deal with it in one way or another. In our implementation, we set the value of the metric to be 0 in these cases for simplicity.
The exact details of how the distribution of recall is derived are given in Algorithm 1. We assume that the input distributions are given in the form of associative arrays, where the values of the random variables act as keys (ranging from 0 to $n_{+}$ or $n_{-}$) and each value corresponding to a given key represents the respective probability. The worst-case time complexity of the algorithm is $\mathcal{O}(n^{2})$ since there are $n_{+}\cdot n_{-}$ iteration steps with $n_{+}+n_{-}=n$ and given a hash table implementation, the search and insertion operations can be performed in $\mathcal{O}(n)$ time, while all other operations are constant time.
Data: $(X_{TP},X_{FN})$ (Distributions for $TP$ and $FN$)
Result: $X_{recall}$   (The full distribution of recall)
$n_{+}\leftarrow|X_{TP}|-1$;
$n_{-}\leftarrow|X_{FN}|-1$;
$X_{recall}\leftarrow\{\}$;    (Empty associative array);
$X_{recall}[0]\leftarrow X_{TP}[0]$;
$X_{recall}[1]\leftarrow X_{FN}[0]-(X_{TP}[0]\cdot X_{FN}[0])$;
for $i\leftarrow 1$ to $n_{+}$ do
       for $j\leftarrow 1$ to $n_{-}$ do
             if key $\frac{i}{i+j}$ already exists then
                   $X_{recall}\left[\frac{i}{i+j}\right]\leftarrow X_{recall}\left[\frac{i}{i+j}\right]+X_{TP}[i]\cdot X_{FN}[j]$;
            else
                   $X_{recall}\left[\frac{i}{i+j}\right]\leftarrow X_{TP}[i]\cdot X_{FN}[j]$;
             end if
       end for
end for
return $X_{recall}$
@@
For F1, we make use of the formula given in Equation (15), and write this in terms of random variables.
$X_{F_{1}}=\frac{2\cdot X_{TP}}{2\cdot X_{TP}+X_{FP}+X_{FN}}$ (8)
and use the fact that $X_{TP}+X_{FP}=n_{+}$, which allows us to rewrite Equation (8) as
$X_{F_{1}}=\frac{2\cdot X_{TP}}{X_{TP}+X_{FN}+n_{+}},$ (9)
where $n_{+}$ is fixed. Similarly to recall, we can form the distribution $X_{F_{1}}$ by iterating over the joint distribution of $X_{TP}$ and $X_{FN}$ and collecting all encountered values of the expression on the right side of Equation (9) with their respective probabilities. The details of how the distribution of F1 is derived are given in Algorithm 2 with the same assumptions and notation as was used with Algorithm 1. The worst-case time complexity of the algorithm is identical to Algorithm 1, namely $\mathcal{O}(n^{2})$.
Data: $(X_{TP},X_{FN})$ (Distributions for $TP$ and $FN$)
Result: $X_{F_{1}}$     (The full distribution of F1)
$n_{+}\leftarrow|X_{TP}|-1$;
$n_{-}\leftarrow|X_{FN}|-1$;
$X_{F_{1}}\leftarrow\{\}$;     (Empty associative list);
$X_{F_{1}}[0]\leftarrow X_{TP}[0]$;
for $i\leftarrow 1$ to $n_{+}$ do
       for $j\leftarrow 0$ to $n_{-}$ do
             if key $X_{F_{1}}\frac{2i}{i+j+n_{+}}$ already exists then
                   $X_{F_{1}}\left[\frac{2i}{i+j+n_{+}}\right]\leftarrow X_{F_{1}}\left[\frac{2i}{i+j+n_{+}}\right]+X_{TP}[i]\cdot X_{FN}[j]$;
            else
                   $X_{F_{1}}\left[\frac{2i}{i+j+n_{+}}\right]\leftarrow X_{TP}[i]\cdot X_{FN}[j]$;
             end if
       end for
end for
return $X_{F_{1}}$
@@ 3.3 Shortcuts for Point estimates
Once the full distribution of a given metric is derived, one can use the expected value of the distribution as a point estimate. Assuming the distribution of a metric of interest contains $m$ different values $x_{i}$ and their respective probabilities $p_{i}$, the expected value is by definition
$\hat{\mu}_{metric}=\sum_{i=1}^{m}x_{i}p_{i}.$ (10)
However, with accuracy and precision, one does not need to form the full distributions of the metrics (or even estimate the confusion matrix) to derive the expectations directly from the confidence scores. We will now show how to apply a shortcut for the expectations of these metrics.
3.3.1 Accuracy
@@
3.3.2 Precision
In a similar fashion to accuracy, we can calculate the expectation directly from the confidence scores by
@@ 3.3.2 Precision
With recall and F1, one can derive highly accurate approximations for the expectations directly from the confidence scores, given that the monitoring window size is sufficiently large. These approximations make use of the following lemma, which we will prove in Appendix B.
@@
Let $X=X_{1}+...+X_{n_{1}}$ and $Y=Y_{1}+...+Y_{n_{2}}$ be random variables following Poisson binomial distributions with $X\perp\!\!\!\!\perp Y$. Moreover, assume that there exist $\delta\in(0,\frac{1}{2}]$ and $N\in\mathbb{N}$ such that for all $n=n_{1}+n_{2}\geq N$, $\delta\leq\frac{n_{1}}{n}\leq 1-\delta$. Finally, assume that $\operatorname{Var}[X]\rightarrow\infty$ and $\operatorname{Var}[Y]\rightarrow\infty$, when $n\rightarrow\infty$ and that $0\leq a\leq 1$. Then,
@@
when $n\to\infty$. Furthermore, the decay rate of the approximation error $\left|\mathbb{E}\left[\frac{X}{X+Y+an}\right]-\frac{\mathbb{E}[X]}{\mathbb{E}[X]+\mathbb{E}[Y]+an}\right|$ is $\mathcal{O}\left(\frac{1}{\sqrt{n}}\right)$.
3.3.3 Recall
@@
$\mathbb{E}[X_{recall}]$ $=\mathbb{E}\left[\frac{X_{TP}}{X_{TP}+X_{FN}}\right]\approx\frac{\mathbb{E}[X_{TP}]}{\mathbb{E}[X_{TP}]+\mathbb{E}[X_{FN}]}$
$=\frac{\sum\limits_{i\in I^{+}}S_{i}}{\sum\limits_{i\in I^{+}}S_{i}+\sum\limits_{i\in I^{-}}S_{i}}=\frac{\sum\limits_{i\in I^{+}}S_{i}}{\sum\limits_{i=1}^{n}S_{i}}.$
3.3.4 F1
Similarly, with F1, we can use Lemma 2 (with $a=\frac{n_{+}}{n}$) to approximate the expectation with
$\mathbb{E}[X_{F_{1}}]$ $=\mathbb{E}\left[\frac{2\cdot X_{TP}}{X_{TP}+X_{FN}+n_{+}}\right]$
$\approx\frac{2\cdot\mathbb{E}[X_{TP}]}{\mathbb{E}[X_{TP}]+\mathbb{E}[X_{FN}]+n_{+}}$
$=\frac{2\cdot\sum\limits_{i\in I^{+}}S_{i}}{\sum\limits_{i\in I^{+}}S_{i}+\sum\limits_{i\in I^{-}}S_{i}+n_{+}}$
$=\frac{2\cdot\sum\limits_{i\in I^{+}}S_{i}}{\sum\limits_{i=1}^{n}S_{i}+n_{+}}.$
@@
3.4 Confidence Bands
As we have seen, the full distributions of our four metrics of interest are not needed to find (or at least approximate) the expected values, which can then be used as a point estimate for any of the metrics. However, having access to the full distributions comes in handy when one wants to take the uncertainty in these estimates into account. This gives a more fine-grained picture of the behavior of a metric of interest under a monitoring setting.
For example, the 95% Highest Density Interval (HDI) is the smallest interval, which is guaranteed to contain the true value of the sample metric with 95% probability. In Appendix C, we construct an algorithm, which is guaranteed to find the $(1-\alpha)$ HDI, where $\alpha$ is a user set parameter ($\alpha=0.05$ corresponds with a 95% confidence interval) with time complexity $\mathcal{O}(n\log n)$.
@@ 4 Experiments
In this section, we conduct some experiments to back up our theoretical claims. We begin by analyzing the empirical convergence rate of the shortcut estimators presented in Section 3.3 and then move to experiments with real-life data. Additional experiments can be found in Appendix D.
@@ 4.1 Convergence of the Shortcut Estimators
We used Lemma (2) to give asymptotic guarantees for the approximations of the expectations of recall and F1. Here we will demonstrate experimentally the rate of convergence in practice. For this, we created synthetic data by sampling confidence scores from different beta distributions, derived the full distributions for the metrics, and calculated the expected values. We then compared these values to the approximations derived using the shortcut formulas presented in Section 3.3 and recorded the approximation error. We repeated this procedure 10,000 times for each monitoring window size, sampling from a different beta distribution, whose $\alpha$ and $\beta$ parameters had been (uniformly) randomly selected from the interval $[0.1,10]$ in each trial. We used monitoring window sizes ranging from 10 to 1,000. The results are shown in Figure 1.
@@
The figure shows that the approximation error decreases rapidly, with the mean absolute error already below 0.001 at window size 100. The deviation of the approximation error also goes to zero quite rapidly, which means that the shortcuts yield consistent estimates of the metrics in question. It is to be noted that this analysis does not depend in any way on whether the confidence scores are calibrated or not. Based on these findings, we recommend deriving the full distributions of the estimated metrics for small monitoring window sizes and using the shortcut estimators when the monitoring window size is big (500+ depending on the situation and user needs) and when one is not interested in the confidence intervals of the estimates.
@@ 4.2 Tableshift Benchmark
For experiments with real-life data, we used datasets from the recently published TableShift benchmark [Gardner et al., 2024]. It provides an API to access 15 tabular datasets collected from various sources to test the robustness of binary classifiers under distribution shift. Here, we use the benchmark outside of its original scope to see how well CBPE is able to estimate different metrics within and outside the original data distribution.
Unfortunately, the API is not well maintained, which made accessing most of the datasets somewhat difficult. For most datasets, we overcame these issues, but had to leave three datasets out of our experiments due to being unable to access them. In addition, four of the datasets require credentialed access, so these were also left out of our experiments. The remaining eight still present a heterogeneous collection of datasets [^3] Full descriptions of these datasets are given at
https://tableshift.org/datasets.html
from diverse sources and domains to offer relevant insight into the applicability of our proposed method.
We trained an XGBoost model for each of the remaining eight datasets. We optimized the hyperparameters using Bayesian optimization over the search spaces defined in the original benchmark. The performance of the models was then estimated with CBPE for all four metrics of interest using unseen test sets from both in-distribution (ID) and out-of-distribution (OoD) splits. We ran 1,000 trials using a monitoring window size of 500. That is, for each trial, we sampled 500 random data points from the test sets. We then extracted the true values of the metrics and calculated the estimation error of the CBPE method. The results are presented in Table 2, where for both splits we report class imbalance (the prevalence of the positive class) as "Bias", calibration error of each model as "ACE", the actual values for the four metrics of interest, and the estimation errors for the CBPE method as Mean Absolute Error (MAE). We also provide error bounds for the estimation errors in the form of two standard deviations. All values are given as percentages.
@@
The results show that CBPE results in low estimation errors when the calibration error is relatively small and in some cases even when the calibration error is rather large as is the case with Hospital Readmission data. It is worthwhile to notice that all of these models are optimized for accuracy and their performance with regards to other metrics might be lacking. This is most notable with the Diabetes data. This shows that CBPE can perform well for a wide range of true metric values. In terms of accuracy, the performance of all models drops significantly with the shift. This is most dramatic for the model trained with ASSISTments data, which also becomes markedly uncalibrated after the shift. Finally, the performance of the models trained for the Hospital Readmission and Hypertension datasets is lacking to begin with. Both are only slightly better than random guessing indicating that the models struggle to find generalizable patterns from the training sets. The Hospital Readmission model is not calibrated even for the ID data, but its calibration error does not increase dramatically with the shift, whereas the Hypertension model is roughly calibrated in the ID split, but highly uncalibrated in the OoD split.
The deviation in the estimation errors is relatively small for accuracy and precision, and somewhat higher for recall and F1, except for the model trained with the Unemployment data, which has notably high variance in all estimation errors except for accuracy. The correlations between calibration errors and estimation errors were very strong for both accuracy and precision ($\rho=0.960$ and $\rho=0.983$ respectively) but interestingly less so for recall and F1 ($\rho=0.199$ and $\rho=0.873$ respectively).

### v2

@@ 4.2.3 Recall
As with accuracy, the joint distribution of $X_{TP}$ and $X_{FN}$ is the product of the marginal distributions since $X_{TP}\perp\!\!\!\!\perp X_{FN}$. Thus, we can form the distribution of $X_{recall}$ by iterating over the joint $P(X_{TP},X_{FN})$ and collecting all encountered values of the expression on the right side of the Equation (5) with their respective probabilities. A small shortcut is to realize that $X_{recall}=0$ whenever $X_{TP}=0$, and $X_{recall}=1$ whenever $X_{FN}=0$, with the exception of both $X_{TP}$ and $X_{FN}$ being zero, which means that recall is undefined. In most practical situations, the probability of this event is negligible, but a user should be prepared to deal with it in one way or another. In our implementation, we set the value of the metric to be 0 in these cases for simplicity.
The exact details of how the distribution of recall is derived are given in Algorithm 1. We assume that the input distributions are given in the form of associative arrays, where the values of the random variables act as keys (ranging from 0 to $n_{+}$ or $n_{-}$) and each value corresponding to a given key represents the probability of observing that key. On lines 1-3 we collect the maximum key value for both $X_{TP}$ and $X_{FN}$ and initialize an empty associative array for the distribution of $X_{recall}$. On line 4, we set the probability of recall being zero to be equal to the probability that the number of true positives is zero. On line 5 we set the probability of recall being one to be equal to the probability that the number of false negatives is zero minus the probability of true positives and false negatives both being zero. On lines 6-14 we iterate over all combinations $(i,j)$ of non-zero keys with $i$ from $X_{TP}$ and $j$ from $X_{FN}$, and check if the key $\frac{i}{i+j}$ already exists in the associative array describing the distribution of $X_{recall}$. If not, the probability of observing key $\frac{i}{i+j}$ will become the product of the probabilities $X_{TP}[i]$ and $X_{FN}[j]$. Otherwise, the probability of observing the key $\frac{i}{i+j}$ is increased by the amount of that product.
The worst-case time complexity of the algorithm is $\mathcal{O}(n^{2})$ since there are $n_{+}\cdot n_{-}$ iteration steps with $n_{+}+n_{-}=n$ and given a hash table implementation, the search and insertion operations can be performed in $\mathcal{O}(n)$ time, while all other operations are constant time [^2] The Poisson binomial distributions for the confusion matrix elements can be constructed using an efficient algorithm based on the Fast Fourier Transform with a time complexity of $\mathcal{O}(n\log n)$ [hong:2013]..
@@
[…]
8:for $i\leftarrow 1$ to $n_{+}$ do
9:  for $j\leftarrow 1$ to $n_{-}$ do
10:   if key $\frac{i}{i+j}$ already exists in $X_{recall}$ then
11:     $X_{recall}\left[\frac{i}{i+j}\right]\leftarrow X_{recall}\left[\frac{i}{i+j}\right]+X_{TP}[i]\cdot X_{FN}[j]$
12:   else
13:     $X_{recall}\left[\frac{i}{i+j}\right]\leftarrow X_{TP}[i]\cdot X_{FN}[j]$
14:   end if
15:  end for
16:end for
@@
F1 can be written as a derived random variable with the expression
$X_{F_{1}}=\frac{2\cdot X_{TP}}{2\cdot X_{TP}+X_{FP}+X_{FN}}=\frac{2\cdot X_{TP}}{X_{TP}+X_{FN}+n_{+}},$ (6)
where $n_{+}$ is fixed. Similarly to recall, we can form the distribution $X_{F_{1}}$ by iterating over the joint distribution of $X_{TP}$ and $X_{FN}$ and collecting all encountered values of the expression on the right side of Equation (6) with their respective probabilities. The details of how the distribution of F1 is derived are given in Algorithm 2 with the same assumptions and notation as was used with Algorithm 1. The logic in Algorithm 2 is completely analogous to that of Algorithm 1. The only difference is in replacing the formula for calculating recall with the formula for calculating F1. The worst-case time complexity of the algorithm is also identical to Algorithm 1, namely $\mathcal{O}(n^{2})$.
@@
[…]
7:for $i\leftarrow 1$ to $n_{+}$ do
8:  for $j\leftarrow 0$ to $n_{-}$ do
9:   if key $\frac{2i}{i+j+n_{+}}$ already exists in $X_{F_{1}}$ then
10:     $X_{F_{1}}\left[\frac{2i}{i+j+n_{+}}\right]\leftarrow X_{F_{1}}\left[\frac{2i}{i+j+n_{+}}\right]+X_{TP}[i]\cdot X_{FN}[j]$
11:   else
12:     $X_{F_{1}}\left[\frac{2i}{i+j+n_{+}}\right]\leftarrow X_{TP}[i]\cdot X_{FN}[j]$
13:   end if
14:  end for
15:end for
@@ 4.3 Shortcuts for Point Estimates
Once the full distribution of a given metric is derived, one can use the expected value of the distribution as a point estimate. However, with accuracy and precision, one does not need to form the full distributions of the metrics (or even estimate the confusion matrix) to derive the expectations directly from the confidence scores. We will now show how to apply a shortcut for the expectations of these metrics.
4.3.1 Accuracy and Precision
@@
Similarly, we can calculate the expectation of precision directly with
@@
4.3.2 Recall and F1
With recall and F1, and a sufficiently large monitoring window, highly accurate approximations of expectations can be calculated using the confidence scores. These approximations rely on the following lemma, which we will prove in the Appendix B.
@@ 4.3.2 Recall and F1
when $n\to\infty$. Furthermore, the decay rate of the approximation error is $\mathcal{O}\left(\frac{1}{\sqrt{n}}\right)$.
@@
$\mathbb{E}[X_{recall}]=\mathbb{E}\left[\frac{X_{TP}}{X_{TP}+X_{FN}}\right]\approx\frac{\mathbb{E}[X_{TP}]}{\mathbb{E}[X_{TP}]+\mathbb{E}[X_{FN}]}=\frac{\sum\limits_{i\in I^{+}}S_{i}}{\sum\limits_{i\in I^{+}}S_{i}+\sum\limits_{i\in I^{-}}S_{i}}=\frac{\sum\limits_{i\in I^{+}}S_{i}}{\sum\limits_{i=1}^{n}S_{i}}.$
Similarly, we can use Lemma 2 (with $a=\frac{n_{+}}{n}$) to approximate the expectation of F1 as
$\mathbb{E}[X_{F_{1}}]=\mathbb{E}\left[\frac{2\cdot X_{TP}}{X_{TP}+X_{FN}+n_{+}}\right]\approx\frac{2\cdot\mathbb{E}[X_{TP}]}{\mathbb{E}[X_{TP}]+\mathbb{E}[X_{FN}]+n_{+}}=\frac{2\cdot\sum\limits_{i\in I^{+}}S_{i}}{\sum\limits_{i=1}^{n}S_{i}+n_{+}}.$
@@ 5 Experiments
In this section, we conduct some experiments to back up our theoretical claims. We begin by analyzing the empirical convergence rate of the shortcut estimators presented in Section 4.3. Then, we examine the quality of the confidence intervals of our proposed estimator, and finally, move to experiments with real-life data. [^3] We make the code available at https://github.com/JuhaniK/CBPE-experiments
@@ 5.1 Convergence of the Shortcut Estimators
We used Lemma (2) to give asymptotic guarantees for the approximations of the expectations of recall and F1. Here, we demonstrate experimentally the rate of convergence in practice. For this, we created synthetic data by sampling confidence scores from different beta distributions, derived the full distributions for the metrics, and calculated the expected values. We then compared these values to the approximations derived using the shortcut formulas presented in Section 4.3 and recorded the approximation error. We repeated this procedure 10,000 times for each monitoring window size, sampling from a different beta distribution, whose $\alpha$ and $\beta$ parameters had been (uniformly) randomly selected from the interval $[0.1,10]$ in each trial. We used monitoring window sizes ranging from 10 to 1,000. The results are shown in Figure 1.
@@
The figure shows that the approximation error decreases rapidly, with the mean absolute error already below 0.001 at window size 100. The deviation of the approximation error also goes to zero quite rapidly, which means that the shortcuts yield consistent estimates for the metrics in question. It is to be noted that this analysis does not depend in any way on whether the confidence scores are calibrated or not. Based on these findings, we recommend deriving the full distributions of the estimated metrics for small monitoring window sizes and using the shortcut estimators when the monitoring window size is big (500+, depending on the situation and user needs) and when one is not interested in the confidence intervals of the estimates.
5.2 Quality of the Confidence Intervals
In this experiment, we used the same synthetic data set as in the experiment described in Section 5.1, but with the addition of using the sampled confidence scores to sample a corresponding label for each confidence score from a Bernoulli distribution using the confidence score as a parameter. This way of reverse sampling ensures that the simulated classifier is calibrated by construction although a small calibration error might persist due to sampling effects. After this, we again conducted 10,000 trials for each monitoring window size, ranging from 100 to 1,000, and recorded whether the actual value was within a predicted 95% or 90% confidence interval. Finally, we calculated the ratio of these successful trials over all trials, which should be close to 0.95 or 0.9, respectively.
@@ 5.3 Tableshift Benchmark
For experiments with real-life data, we used datasets from the recently published TableShift benchmark [gardner:2024]. It provides an API to access 15 tabular datasets collected from various sources to test the robustness of binary classifiers under distribution shift. Here we use the benchmark outside of its original scope to see how well CBPE estimates different metrics within and outside the original data distribution. Unfortunately, the API is not well maintained, which made accessing most of the datasets somewhat difficult. For most datasets, we overcame these issues, but had to leave three datasets out of our experiments due to being unable to access them. In addition, four of the datasets require credentialed access, so these were also left out of our experiments. The remaining eight still present a heterogeneous collection of datasets [^4] Full descriptions of these datasets are given at https://tableshift.org/datasets.html from diverse sources and domains to offer relevant insight into the applicability of our proposed method.
We trained an XGBoost model for each of the remaining eight datasets. We optimized the hyperparameters using Bayesian optimization over the search spaces defined in the original benchmark. The performance of the models was then estimated with CBPE for all four metrics of interest using unseen test sets from both in-distribution (ID) and out-of-distribution (OoD) splits. We ran 1,000 trials using a monitoring window size of 500. That is, for each trial, we sampled 500 random data points from the test sets. We then extracted the true values of the metrics and calculated the estimation error of the CBPE method. The results are presented in Table 2, where for both splits we report class imbalance (the prevalence of the positive class) as ”Bias”, the Adaptive Expected Calibration Error [nixon:2019] of each model as ”ACE”, the actual values for the four metrics of interest, and the estimation errors for the CBPE method as Mean Absolute Error (MAE). We also provide error bounds for the estimation errors in the form of two standard deviations. All values are given as percentage units.
@@
The results show that CBPE results in low estimation errors when the calibration error is relatively small and in some cases even when the calibration error is rather large, as is the case with Hospital Readmission data. It is worthwhile to notice that all of these models are optimized for accuracy, and their performance with respect to other metrics might be lacking. This is most notable with the Diabetes data. This shows that CBPE can perform well for a wide range of true metric values. In terms of accuracy, the performance of all models drops significantly with the shift. This is most dramatic for the model trained with ASSISTments data, which also becomes markedly uncalibrated after the shift. Finally, the performance of the models trained for the Hospital Readmission and Hypertension datasets is lacking to begin with. Both are only slightly better than random guessing, indicating that the models struggle to find generalizable patterns from the training sets. The Hospital Readmission model is not calibrated even for the ID data, but its calibration error does not increase dramatically with the shift, whereas the Hypertension model is roughly calibrated in the ID split, but highly uncalibrated in the OoD split.
The deviation in the estimation errors is relatively small for accuracy and precision, and somewhat higher for recall and F1, except for the model trained with the Unemployment data, which has notably high variance in all estimation errors except for accuracy. The correlations between calibration errors and estimation errors were very strong for both accuracy and precision ($\rho=0.960$ and $\rho=0.983$, respectively), but interestingly less so for recall and F1 ($\rho=0.199$ and $\rho=0.873$, respectively).